### v1

# XMAD-Bench: Cross-Domain Multilingual Audio Deepfake Benchmark

Abstract.
Recent advances in audio generation led to an increasing number of deepfakes, making the general public more vulnerable to financial scams, identity theft, and misinformation. Audio deepfake detectors promise to alleviate this issue, with many recent studies reporting accuracy rates close to . However, these methods are typically tested in an in-domain setup, where the deepfake samples from the training and test sets are produced by the same generative models. To this end, we introduce XMAD-Bench, a large-scale cross-domain multilingual audio deepfake benchmark comprising 668.8 hours of real and deepfake speech. In our novel dataset, the speakers, the generative methods, and the real audio sources are distinct across training and test splits. This leads to a challenging cross-domain evaluation setup, where audio deepfake detectors can be tested “in the wild”. Our in-domain and cross-domain experiments indicate a clear disparity between the in-domain performance of deepfake detectors, which is usually as high as , and the cross-domain performance of the same models, which is sometimes similar to random chance. Our benchmark highlights the need for the development of robust audio deepfake detectors, which maintain their generalization capacity across different languages, speakers, generative methods, and data sources. Our benchmark is publicly released at https://github.com/ristea/xmad-bench/.
A pie chart that illustrates the distribution of real and fake speech for each language. XMAD-Bench comprises 668.8 hours of real and fake speech samples across seven languages: Arabic (Ar), English (En), German (De), Mandarin Chinese (Zh), Romanian (Ro), Russian (Ru), and Spanish (Es). For each language, there are two sources of real samples, enabling us to organize the dataset in a cross-domain format.
1. Introduction
The recent development of powerful audio generation models, capable of synthesizing realistic speech from text (Casanova et al., 2022; Huang et al., 2022; Ju et al., 2024; Shen et al., 2024; Tan et al., 2024) and precisely reproducing voices (Jiang et al., 2023; Lee et al., 2023; Chen et al., 2025), opened the path to new application domains. Unfortunately, these advancements also led to an increase in misuses, especially related to deepfake generation. Indeed, it was recently reported that the number of frauds based on deepfake increased 10-fold from 2022 to 2023111Sumsub Expert Roundtable: The Top KYC Trends Coming in 2024. This is particularly worrying for the audio domain, since it was found that roughly of the general public is not able to tell if a voice is real or fake222Artificial Imposters–Cybercriminals Turn to AI Voice Cloning for a New Breed of Scam. In this context, accurately detecting deepfake audio content is of utter importance.
To date, considerable research efforts have been dedicated to advance audio deepfake detection, most of the recent approaches being based on deep learning models (Chen et al., 2023; Jung et al., 2022; Liu et al., 2023; Tak et al., 2022). Impressively, such models reached or even surpassed the threshold in terms of audio deepfake detection performance on existing benchmarks (Croitoru et al., 2024), such as ASVspoof 2019-LA (Wang et al., 2020) and ASVspoof 2021-LA (Yamagishi et al., 2021). However, due to the limitations of existing audio deepfake datasets, audio deepfake detectors are generally tested in an in-domain setup, where the deepfake samples from the training and test sets are generated by the same audio generation tools. Hence, the reported accuracy levels of current audio deepfake detectors do not reflect the actual performance of these models when tested “in the wild”, where the speaker identity or the generative method remains unknown.
To this end, we introduce a novel benchmark for cross-domain audio deepfake detection in multiple languages. XMAD-Bench, which stands for Cross-Domain Multilingual Audio Deepfake Benchmark, contains 668.8 hours of real and fake speech across seven languages: Arabic, English, German, Mandarin Chinese, Romanian, Russian, and Spanish (see Figure 1). XMAD-Bench comprises a variety of languages, including both widely-spoken (English, Arabic) as well as less popular (Romanian) languages. The dataset is balanced in terms of the real versus fake class distribution, containing 207K real samples and 207K deepfake samples. XMAD-Bench also provides an official three-way split of the data samples into a training set, an in-domain test set and a cross-domain test set, such that speakers are distinct across splits. More importantly, the cross-domain test set contains real audio samples from data sources that are distinct from the training set, and deepfake audio samples generated by a different set of generative methods than the training set. This leads to a challenging cross-domain evaluation setup, which allows audio deepfake detectors to be tested “in the wild”. Due to the fast pace of AI research, we emphasize that generative methods can become obsolete in 3-4 years, so the ratio of known models in the test set can drastically decrease over time. To take into account the passage of time, we refer to the “cross-domain” setting as “in the wild”. Nevertheless, since we provide both in-domain and cross-domain test sets, interested parties can easily combine samples from the two test sets to obtain an “in the wild” setup with a desired ratio.
We conduct experiments with both convolutional and transformer architectures, namely ResNet-18 (He et al., 2016), ResNet-50 (He et al., 2016), AST (Gong et al., 2021), SepTr (Ristea et al., 2022), and wav2vec 2.0 (Baevski et al., 2020). The objective of our experiments is to compare the in-domain and cross-domain performance of neural models with various configurations. Our results show that state-of-the-art models are capable of reaching extremely high accuracy rates (usually close to 100%) on the in-domain test split, but they fail to maintain their strong performance in the cross-domain setting. The obvious disparity between the in-domain and cross-domain performance of state-of-the-art models indicates that more research efforts need to be dedicated to the development of robust audio deepfake detectors, which maintain their generalization capacity across different speakers, generative methods, and data sources.
In summary, our contribution is twofold:
We introduce XMAD-Bench, a large-scale cross-domain multilingual audio deepfake benchmark comprising 668.8 hours of real and deepfake speech across seven languages.
We carry out comprehensive in-domain and cross-domain experiments to evaluate audio deepfake detectors based on state-of-the-art neural architectures, showing that such models exhibit generally poor generalization capacity.
2. Related Work
The scientific community uses a relatively small number of existing datasets to assess the effectiveness of audio deepfake detection methods. Such datasets usually contain a single language, which limits their usage in multilingual scenarios, with only a few exceptions, such as MLAAD (Müller et al., 2024) and WaveFake (Frank and Schönherr, 2021), comprising multiple languages for spoofed audio detection. Some of the most popular datasets correspond to the ASVspoof challenges, especially the 2019 and 2021 editions (Wang et al., 2020; Yamagishi et al., 2021), which encouraged research in anti-spoofing methods for Automatic Speaker Verification (ASV). Both datasets contain only English samples based on the Voice Cloning Toolkit corpus (Yamagishi et al., 2017). ADD 2022 (Yi et al., 2022) and ADD 2023 (Yi et al., 2023) introduce various scenarios, such as low-quality fake audio detection, partially fake audio detection and deepfake algorithm recognition, with a corresponding dataset for each task. They are based on the AISHELL Mandarin speech corpus (Bu et al., 2017; Shi et al., 2021; Fu et al., 2021) and contain fake samples generated with various text-to-speech (TTS) and voice conversion (VC) systems. WaveFake (Frank and Schönherr, 2021) consists of English and Japanese fake samples generated with different TTS models, starting from real clips from the LJSPEECH (Ito and Johnson, 2017) and JSUT (Sonobe et al., 2017) datasets. Reimao et al. (Reimao and Tzerpos, 2019) proposed the Fake or Real (FoR) dataset, with real English clips collected from open-source datasets, such as LJSPEECH (Ito and Johnson, 2017), Arctic, VoxForge, and social media platforms such as YouTube. The fake samples are generated using both open-source and commercial TTS systems.
| Language | Subset | Data source | Fake generation methods | Total | #samples | Average | #speakers |
| length (h) | length (s) | ||||||
| Arabic | Training | CommonVoice | fairseq+KNN-VC, fairseq+FreeVC | 68.46 | 56,114 | 4.39 | 204 |
| In-Domain Test | CommonVoice | fairseq+KNN-VC, fairseq+FreeVC | 24.63 | 20,488 | 4.32 | 52 | |
| Cross-Domain Test | MASC | XTTSv2 | 23.80 | 12,984 | 6.59 | 1,502 | |
| English | Training | CommonVoice | VITS+KNN-VC, XTTSv2 | 114.05 | 75,000 | 5.47 | 885 |
| In-Domain Test | CommonVoice | VITS+KNN-VC, XTTSv2 | 33.99 | 21,368 | 5.73 | 222 | |
| Cross-Domain Test | M-AILABS | VITS+OpenVoice, GlowTTS+FreeVC, | 82.41 | 39,690 | 7.47 | 3 | |
| VALL-E-X, YourTTS | |||||||
| German | Training | CommonVoice | fairseq+OpenVoice, XTTSv2 | 17.73 | 11,672 | 5.47 | 60 |
| In-Domain Test | CommonVoice | fairseq+OpenVoice, XTTSv2 | 4.89 | 3,262 | 5.40 | 15 | |
| Cross-Domain Test | M-AILABS | Tacotron2-DDC+FreeVC, Tacotron2-DDC+KNN-VC, | 6.78 | 3,100 | 7.88 | 5 | |
| VITS+FreeVC, VITS+KNN-VC, YourTTS | |||||||
| Mandarin | Training | CommonVoice | Tacotron2-DDC-GST+KNN-VC, Bark+FreeVC | 16.56 | 11,766 | 5.06 | 821 |
| In-Domain Test | CommonVoice | Tacotron2-DDC-GST+KNN-VC, Bark+FreeVC | 2.81 | 2,010 | 5.04 | 45 | |
| Cross-Domain Test | AISHELL-3 | MeloTTS+OpenVoice, | 1.85 | 2,002 | 3.33 | 95 | |
| VALL-E-X, XTTSv2 | |||||||
| Romanian | Training | CommonVoice | VITS+KNN-VC, VITS+FreeVC | 25.20 | 25,934 | 3.50 | 144 |
| In-Domain Test | CommonVoice | VITS+KNN-VC, VITS+FreeVC | 4.72 | 4,886 | 3.47 | 26 | |
| Cross-Domain Test | VoxPopuli | VITS+OpenVoice | 18.77 | 6,672 | 10.13 | 38 | |
| Russian | Training | CommonVoice | VITS+KNN-VC, XTTSv2 | 86.65 | 56,126 | 5.56 | 158 |
| In-Domain Test | CommonVoice | VITS+KNN-VC, XTTSv2 | 17.08 | 11,318 | 5.43 | 40 | |
| Cross-Domain Test | M-AILABS | VITS+OpenVoice | 89.52 | 34,702 | 9.29 | 3 | |
| Spanish | Training | CommonVoice | fairseq+OpenVoice, XTTSv2 | 18.75 | 10,436 | 6.47 | 65 |
| In-Domain Test | CommonVoice | fairseq+OpenVoice, XTTSv2 | 4.05 | 2,258 | 6.47 | 17 | |
| Cross-Domain Test | M-AILABS | MeloTTS+FreeVC, MeloTTS+KNN-VC, | 6.13 | 3,070 | 7.19 | 3 | |
| Tacotron2-DDC+FreeVC, Tacotron2-DDC+KNN-VC, | |||||||
| VITS+FreeVC, VITS+KNN-VC, YourTTS | |||||||
| Overall | Training | - | - | 347.40 | 247,048 | 5.21 | 2,337 |
| In-Domain Test | - | - | 92.17 | 65,590 | 5.13 | 417 | |
| Cross-Domain Test | - | - | 229.26 | 102,220 | 8.23 | 1,649 |
To evaluate the performance of deepfake detection models in a cross-dataset scenario, Müller et al. (Müller et al., 2022) trained several detectors on the ASVspoof2019 (Wang et al., 2020) dataset and reported results on their novel dataset, called MLAAD. They observed a large performance drop, questioning the generalization capability of various deepfake detectors. MLAAD (Müller et al., 2024) includes samples synthesized from the M-AILABS Speech Dataset (Solak, 2019), being one of the few datasets that contains fake audio samples covering multiple languages. For the majority of languages (not part of M-AILABS), the authors generated fake samples by translating English text samples into additional target languages and then using state-of-the-art TTS models. Despite the large number of languages covered by MLAAD, it contains only fake samples. Therefore, to train a deepfake detector, one also needs a dataset of real samples, containing exactly the same set of languages as MLAAD, but this is not the case for M-AILABS. If the training language sets for real and fake samples are different, the deepfake detector can suffer from significant language biases, i.e. it can mislabel samples in languages for which the real or fake samples are missing. In contrast, we mitigate this problem by constructing a balanced dataset that includes both real and synthesized samples in all target languages.
Li et al. (Li et al., 2024) identified the generalization issue of deepfake detectors, proposing a monolingual cross-domain dataset for audio deepfake detection, called CD-ADD. The dataset contains 300 hours of speech generated by five zero-shot TTS models, hence the cross-domain nature. The authors also make use of the ASVSpoof2019 dataset, as well as pre-trained speech encoders, such as wav2vec 2.0 (Baevski et al., 2020) and Whisper (Radford et al., 2023). The dataset is affected by several perturbations, called “attacks”, which simulate real-world noise and significantly affect the detection performance. The work does not use any real speech data and does not make an explicit effort to accommodate multiple languages.
The results obtained by the latest deepfake detection models tend to saturate existing benchmarks, with GNN-based methods, like AASIST (Jung et al., 2022; Tak et al., 2022), transformer-based methods, like Rawformer (Liu et al., 2023), and other models (Roselló Casado et al., 2023; Truong et al., 2024) reporting EER values of around 1%. We conjecture that the performance of such models would drop significantly when these models are tested on a dataset that would include audio clips from diverse speakers, recorded in different conditions and with spoofed samples generated by different methods. To the best of our knowledge, XMAD-Bench is the first multilingual cross-domain dataset for audio deepfake detection, containing both real and fake samples. Moreover, XMAD-Bench is the largest dataset of its kind, being more than twice as large as recent large-scale datasets, such as MLAAD (Müller et al., 2024) and CD-ADD (Li et al., 2024).
3. Dataset
Overview.
The XMAD-Bench dataset consists of real audio clips collected from various existing speech datasets, and corresponding fake clips generated based on the real ones. Each real clip has a matching fake version, generated using text-to-speech (TTS) and voice conversion (VC) tools, which preserve the text content and speech characteristics of the original. The dataset includes speech samples in Arabic, English, German, Mandarin Chinese, Romanian, Russian, and Spanish. For each language, the clips are sourced from two datasets, one designated for training and in-domain testing, and another for cross-domain testing. The in-domain data is divided into a training split and an in-domain test split. The fake samples for the in-domain data are generated by two distinct generative methods, such that half of the fake audio files are produced by the first method, and the other half by the second method. For the cross-domain test set, fake samples are synthesized by multiple generative methods, distinct from the first two. Moreover, the speakers are distinct across all three splits.
In Table 1, we list the sources of real samples and the generative methods used for each language. For the in-domain data, the real speech samples are collected from Common Voice (Ardila et al., 2020), a massively-multilingual speech corpus. The real samples for the cross-domain test are gathered from the Massive Arabic Speech Corpus (MASC) (Al-Fetyani et al., 2023) (for Arabic), the M-AILABS dataset (Solak, 2019) (for English, German, Russian and Spanish), the AISHELL-3 corpus (Shi et al., 2021) (for Mandarin Chinese), and the VoxPopuli dataset (Wang et al., 2021) (for Romanian). MASC contains audio collected from YouTube in Arabic. AISHELL-3 is a corpus comprising 85 hours of speech from 218 Mandarin speakers. M-AILABS is a multilingual corpus based on audiobooks, while VoxPopuli is a multilingual speech corpus consisting of European Parliament recordings. The datasets are chosen due to their permissive license agreements, which allow us to share data and derivatives for non-commercial research purposes.
A figure presenting the general flow for fake sample generation based on various text-to-speech and voice conversion tools.
Synthesis procedure.
For fake audio generation, the transcribed text of the real audio sample is first passed to a TTS model. Next, a VC tool is applied to the generated speech, using the voice of the speaker uttering the real sample as reference. The general flow used to generate fake audio samples is illustrated in Figure 2. For the TTS step, we alternate between various models, namely VITS (Kim et al., 2021), XTTSv2 (Casanova et al., 2024), fairseq (based on VITS) (Pratap et al., 2024), Tacotron2 (Shen et al., 2018), MeloTTS (Zhao et al., 2023), YourTTS (Casanova et al., 2022), GlowTTS (Kim et al., 2020), VALL-E-X (Zhang et al., 2023) and Bark (Suno, 2023). For the VC step, we alternatively employ KNN-VC (Baas et al., 2023), FreeVC (Li et al., 2022) and OpenVoice (Qin et al., 2024). The TTS and VC models are chosen based on their public availability, state-of-the-art performance and support for the various languages included in XMAD-Bench. Most models, e.g. VITS, XTTSv2, fairseq, Tacotron2, YourTTS, GlowTTS and Bark, are imported from the Coqui TTS library333https://github.com/idiap/coqui-ai-TTS. Other models, e.g. KNN-VC, VALL-E-X, MeloTTS, are imported from the official repositories of the corresponding papers. While some models, such as KNN-VC, FreeVC, and OpenVoice, use the clips synthesized by a TTS model as input, XTTSv2, YourTTS and VALL-E-X take both the transcript and the reference voice, and directly generate fake samples, requiring no additional VC tool. All models are pre-trained on the target languages, requiring no adaptation from our end. Note that fairseq models exhibit lower performance than the other ones, but we decided to keep them due to their multilingual support.
We employ two alternative synthesis procedures per language for the in-domain data, aiming to increase the variability of the fake samples. The cross-domain test set is generated with as many methods per language as possible, all of them being distinct from those used for the in-domain data. All the methods extract the speech characteristics from the real samples that are cloned.
For all databases, we employ random sampling to reach the speech lengths reported in Table 1. For VoxPopuli, in particular, we select clips shorter than 20 seconds, in order to limit variation in clip duration, and discard speakers with less than two minutes of total content duration. The latter threshold is also used to enhance conversion quality when employing KNN-VC. With the exception of KNN-VC, the employed VC tools require a single reference clip. For KNN-VC, multiple clips totaling two minutes are used as reference speech, which improves its VC quality.
| Real | Fake | |
| SAR | 0.8563 | 0.8894 |
| SNR | 34.54 | 34.94 |
The MASC dataset contains many clips from YouTube that vary in length, and may even include music. As music content can be a bias for real clips, we filter clips with a percentage of music content greater than 5%, using a music detection tool from the TVSM (Hung et al., 2022) dataset repository. As MASC is a massive corpus, only the first 7 minutes from each clip are taken and then split into smaller segments, each corresponding to a single caption. After segmentation, only clips with a duration greater than 4 seconds and a caption shorter than 160 characters are kept, to achieve constancy in clip duration, in concordance with the other datasets. MASC does not provide speaker identities, so we simply assume that each speaker appears in only one video. This does not affect the cross-domain nature of XMAD-Bench, since the speakers in CommonVoice and MASC are different.
| Language | Method | In-Domain | Cross-Domain | ||||
| ACC | AUC | EER | ACC | AUC | EER | ||
| Arabic | ResNet-18 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 37.80 (10.11) | 41.26 (12.02) | 57.26 (10.29) |
| ResNet-50 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 25.09 ( 4.03) | 29.67 ( 2.37) | 68.50 ( 2.34) | |
| SepTr | 96.84 ( 1.32) | 98.66 ( 1.09) | 5.04 ( 2.61) | 36.15 ( 8.44) | 24.88 ( 6.32) | 60.56 ( 4.96) | |
| AST | 99.97 ( 0.01) | 99.99 ( 0.00) | 0.01 ( 0.01) | 73.39 ( 1.54) | 81.04 ( 0.02) | 26.68 ( 0.23) | |
| wav2vec 2.0 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 75.03 ( 2.11) | 83.27 ( 0.18) | 25.81 ( 0.54) | |
| English | ResNet-18 | 100.0 ( 0.00) | 99.99 ( 0.01) | 0.00 ( 0.00) | 47.76 ( 3.01) | 47.34 ( 8.84) | 56.63 ( 7.99) |
| ResNet-50 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 54.02 ( 9.64) | 59.56 (11.94) | 45.53 ( 9.31) | |
| SepTr | 93.21 ( 0.79) | 94.67 ( 2.73) | 12.29 ( 3.87) | 42.30 ( 8.66) | 39.20 (16.01) | 59.52 (12.04) | |
| AST | 99.20 ( 0.32) | 99.96 ( 0.02) | 0.66 ( 0.17) | 69.19 ( 1.48) | 79.70 ( 1.78) | 28.52 ( 1.45) | |
| wav2vec 2.0 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 69.67 ( 3.22) | 80.03 ( 1.72) | 28.77 ( 1.55) | |
| German | ResNet-18 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 90.31 ( 6.08) | 99.49 ( 0.34) | 3.65 ( 1.54) |
| ResNet-50 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 96.96 ( 1.09) | 99.79 ( 0.02) | 2.16 ( 0.15) | |
| SepTr | 99.95 ( 0.43) | 99.25 ( 0.22) | 3.82 ( 0.71) | 79.62 ( 1.28) | 86.54 ( 2.19) | 19.95 ( 1.32) | |
| AST | 99.65 ( 0.09) | 99.99 ( 0.00) | 0.32 ( 0.07) | 91.31 ( 2.40) | 97.50 ( 1.03) | 8.68 ( 2.27) | |
| wav2vec 2.0 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 94.94 ( 2.03) | 98.29 ( 1.65) | 3.72 ( 0.67) | |
| Mandarin | ResNet-18 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 68.79 ( 3.09) | 87.08 ( 3.16) | 21.30 ( 3.63) |
| ResNet-50 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 67.31 ( 2.81) | 87.55 ( 1.39) | 20.50 ( 1.63) | |
| SepTr | 96.14 ( 1.44) | 99.07 ( 0.83) | 4.30 ( 2.35) | 65.10 ( 4.83) | 78.82 ( 2.59) | 28.50 ( 2.60) | |
| AST | 99.28 ( 0.20) | 99.95 ( 0.02) | 0.82 ( 0.28) | 77.52 ( 1.10) | 84.77 ( 0.66) | 23.20 ( 1.02) | |
| wav2vec 2.0 | 99.97 ( 0.01) | 99.99 ( 0.00) | 0.01 ( 0.01) | 71.04 ( 2.71) | 83.78 ( 2.01) | 25.43 ( 1.98) | |
| Romanian | ResNet-18 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 61.59 ( 4.96) | 96.32 ( 1.37) | 8.53 ( 2.32) |
| ResNet-50 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 76.39 (10.48) | 93.16 ( 2.04) | 13.83 ( 2.30) | |
| SepTr | 98.51 ( 1.61) | 98.55 ( 2.00) | 3.96 ( 4.62) | 48.31 ( 1.56) | 45.98 (16.68) | 52.55 (12.43) | |
| AST | 99.32 ( 0.39) | 99.98 ( 0.01) | 0.43 ( 0.21) | 57.18 ( 3.76) | 92.60 ( 2.07) | 13.25 ( 3.22) | |
| wav2vec 2.0 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 78.12 ( 3.21) | 93.21 ( 1.99) | 12.09 ( 1.39) | |
| Russian | ResNet-18 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 47.67 ( 3.72) | 36.70 ( 1.99) | 60.38 ( 1.25) |
| ResNet-50 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 33.85 (11.35) | 25.99 (16.72) | 69.11 (12.77) | |
| SepTr | 94.96 ( 1.08) | 82.29 (22.36) | 20.83 (20.38) | 29.22 ( 3.36) | 52.35 (20.57) | 49.43 (15.45) | |
| AST | 99.69 ( 0.16) | 99.99 ( 0.05) | 0.30 ( 0.15) | 66.24 ( 2.67) | 74.07 ( 3.92) | 32.26 ( 3.65) | |
| wav2vec 2.0 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 61.79 ( 3.92) | 68.54 ( 2.03) | 35.66 ( 3.78) | |
| Spanish | ResNet-18 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 90.21 ( 1.68) | 98.78 ( 0.56) | 4.90 ( 1.19) |
| ResNet-50 | 100.0 ( 0.00) | 100.0 ( 0.00) | 0.00 ( 0.00) | 89.72 ( 2.74) | 99.27 ( 0.22) | 4.07 ( 0.58) | |
| SepTr | 96.47 ( 0.59) | 98.98 ( 0.61) | 4.78 ( 1.99) | 75.22 ( 8.58) | 79.19 (14.33) | 24.52 (12.22) | |
| AST | 98.80 ( 0.28) | 99.95 ( 0.03) | 0.97 ( 0.33) | 93.77 ( 1.60) | 98.68 ( 0.59) | 5.73 ( 1.70) | |
| wav2vec 2.0 | 99.97 ( 0.01) | 99.99 ( 0.00) | 0.01 ( 0.01) | 94.62 ( 1.44) | 98.87 ( 0.38) | 4.76 ( 0.78) | |
| Cross-lingual | ResNet-18 | 99.98 ( 0.01) | 99.99 ( 0.00) | 0.01 ( 0.01) | 87.87 ( 0.99) | 94.41 ( 1.37) | 12.52 ( 1.32) |
| ResNet-50 | 99.96 ( 0.01) | 99.99 ( 0.00) | 0.02 ( 0.01) | 86.10 ( 2.64) | 93.51 ( 1.30) | 13.67 ( 1.78) | |
| SepTr | 91.70 ( 4.26) | 82.76 (15.66) | 22.05 (15.92) | 72.67 ( 4.22) | 55.87 (13.06) | 45.69 (10.24) | |
| AST | 98.53 ( 0.46) | 99.87 ( 0.06) | 1.35 ( 0.42) | 81.61 ( 0.70) | 89.77 ( 1.36) | 16.91 ( 0.91) | |
| wav2vec 2.0 | 99.97 ( 0.01) | 99.99 ( 0.00) | 0.01 ( 0.01) | 89.07 ( 0.56) | 93.58 ( 1.16) | 11.82 ( 1.07) | |
Finally, all real and fake clips are trimmed for silence at the start and the end, and resampled to 16 kHz. This ensures that real and fake samples cannot be distinguished based on the sampling rate or the presence of silence periods.
Statistics.
XMAD-Bench is composed of 668.8 hours of real and fake content coming from 4,403 different speakers. As shown in Table 1, there is a noticeable variation in average clip duration and speaker diversity, across domains and source datasets. Since the M-AILABS dataset is composed of audiobooks, its speaker diversity is quite low. The average length of audio clips in VoxPopuli is generally longer, even after we discarded clips longer than 20 seconds. However, these differences do not introduce any spurious correlations in the dataset, since the training split is consistent across languages.
To assess the quality differences between real and fake samples, we report the speech activity ratio (SAR) and the signal-to-noise ratio (SNR) in Table 2. We highlight that the SAR and SNR values are fairly similar for both real and fake samples, confirming that the fake samples are of high quality.
4. Experiments
Audio deepfake detectors.
For deepfake detection, we consider four pre-trained models, namely ResNet-18 and ResNet-50 (He et al., 2016) from Torchvision444https://github.com/pytorch/vision, Audio Spectrogram Transformer (AST) (Gong et al., 2021) from Hugging Face555https://huggingface.co/MIT/ast-finetuned-audioset-10-10-0.4593, and wav2vec 2.0 (Baevski et al., 2020; Tak et al., 2022) from Hugging Face666https://huggingface.co/docs/transformers/model_doc/wav2vec2. We fine-tune these models on XMAD-Bench. We also train a fifth model from scratch, namely SepTr (Ristea et al., 2022). The SepTr model777https://github.com/ristea/septr is composed of 3 separable transformer blocks, each with 5 attention heads. The dimension of each head and the dimension of the MLP layer are set to 256. Each audio clip is converted into a spectrogram to be processed by each of the five models.
Experimental setup.
We conduct experiments by training detectors on the training set of each language, and evaluating them on the in-domain test set, after every epoch. The checkpoints achieving the highest performance on the in-domain split are further tested on the cross-domain test set. In addition, we also carry out cross-lingual experiments, training the models on Arabic, German, Romanian, Russian and Spanish, and testing them on English and Mandarin. In the cross-lingual setup, we randomly select at most 3,000 samples per language. In all experiments, each clip is augmented during training with a probability of . Augmentations include time shifting by rolling the signal, speed augmentation, volume augmentation by applying random gain, clipping, reverberation, spectral shifting (high-shelf, low-shelf and peak filters) and pitch shifting.
Hyperparameters.
We optimize all models via the cross-entropy loss. Each model is trained for 20 epochs, using a learning rate of and no weight decay. The mini-batch size for each model depends on the size of the respective model. We thus set the mini-batch size to 200 for ResNet-18, 120 for ResNet-50, 16 for wav2vec 2.0, and 10 for both AST and SepTr. The input length of all models is fixed to 5 seconds. We randomly select a segment of 5 seconds from longer clips, while zero-padding the shorter ones. The spectrograms are generated using a 320-point Short-Time Fourier Transform, with 160 hops in the time-domain, on which we apply a Hann sliding window. The resulting size of a spectrogram is . For efficiency reasons, we downsample the spectrogram by a factor of 4 for SepTr. Unlike the other models, AST is based on Mel spectrograms of components, this being the default configuration for AST. The wav2vec 2.0 model directly consumes raw audio waveforms sampled at 16 kHz, requiring no handcrafted spectral preprocessing.
A figure showing the cross-domain confusion matrices of ResNet-18 (first row) and AST (second row) on Arabic (first column) and English (second column).
Results.
As shown in Table 3, several models reach an impressive performance of 100% on the in-domain split, confirming that deep neural models can easily capture the characteristics of generative models, regardless of the target language. However, the cross-domain results support our conjecture, demonstrating that most metrics drastically decrease when detectors are tested in the cross-domain scenario, where fake samples are generated by models unknown to the detectors. Notably, wav2vec 2.0 outperforms all other models in the cross-domain setting on four of the seven target languages, showing its strong ability to generalize to unseen generative models. Furthermore, wav2vec 2.0 exhibits robust multilingual performance, ranking among the top performers in the multilingual setting and demonstrating its effectiveness in capturing language-independent representations of synthetic speech. The cross-lingual experiments indicate that the language shift does not significantly affect performance.
In Figure 3, we present the confusion matrices for the cross-domain evaluation of ResNet-18 and AST on Arabic and English, respectively. On Arabic, AST tends to label many real samples as fake, while ResNet-18 exhibits both types of mistakes. On English, the two models have opposite biases. Overall, the confusion matrices indicate that the biases are specific to the models, not to XMAD-Bench. This observation further confirms that XMAD-Bench is a challenging and robust benchmark.
5. Conclusion and Future Work
In this paper, we introduced a novel multilingual cross-domain audio dataset to evaluate deepfake detectors “in the wild”. We discussed the methodology used to generate diverse fake clips for three partitions (training, in-domain test, and cross-domain test), and we further presented dataset statistics across the seven target languages. We evaluated the performance of five state-of-the-art models from the literature in terms of multiple metrics, showing that our cross-domain evaluation scenario causes a significant decline for all the reported metrics. While we were able to replicate the near perfect in-domain performance previously reported in literature, our cross-domain results highlighted the difficulty of performing audio deepfake detection across datasets, languages, speakers, and deepfake generative methods.
In future work, we will focus on the development of robust domain adaptation techniques to improve the results in the cross-domain setup, which simulates a challenging real-world scenario.
Acknowledgements.
This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CCCDI - UEFISCDI, project number PN-IV-P6-6.3-SOL-2024-2-0227, within PNCDI IV.References
- (1)
- Al-Fetyani et al. (2023) Mohammad Al-Fetyani, Muhammad Al-Barham, Gheith Abandah, Adham Alsharkawi, and Maha Dawas. 2023. MASC: Massive Arabic Speech Corpus. In SLT. 1006–1013.
- Ardila et al. (2020) R. Ardila, M. Branson, K. Davis, M. Henretty, M. Kohler, J. Meyer, R. Morais, L. Saunders, F. M. Tyers, and G. Weber. 2020. Common Voice: A Massively-Multilingual Speech Corpus. In LREC. 4218–4222.
- Baas et al. (2023) Matthew Baas, Benjamin van Niekerk, and Herman Kamper. 2023. Voice Conversion With Just Nearest Neighbors. In INTERSPEECH. 2053–2057.
- Baevski et al. (2020) Alexei Baevski, Yuhao Zhou, Abdelrahman Mohamed, and Michael Auli. 2020. wav2vec 2.0: A framework for self-supervised learning of speech representations. In NeurIPS, Vol. 33. 12449–12460.
- Bu et al. (2017) Hui Bu, Jiayu Du, Xingyu Na, Bengu Wu, and Hao Zheng. 2017. AISHELL-1: An Open-Source Mandarin Speech Corpus and A Speech Recognition Baseline. In O-COCOSDA.
- Casanova et al. (2024) Edresson Casanova, Kelly Davis, Eren Gölge, Görkem Göknar, Iulian Gulea, Logan Hart, Aya Aljafari, Joshua Meyer, Reuben Morais, Samuel Olayemi, and Julian Weber. 2024. XTTS: a Massively Multilingual Zero-Shot Text-to-Speech Model. In INTERSPEECH. 4978–4982.
- Casanova et al. (2022) Edresson Casanova, Julian Weber, Christopher D. Shulby, Arnaldo Candido Junior, Eren Gölge, and Moacir A. Ponti. 2022. YourTTS: Towards Zero-Shot Multi-Speaker TTS and Zero-Shot Voice Conversion for Everyone. In ICML. 2709–2720.
- Chen et al. (2023) Feng Chen, Shiwen Deng, Tieran Zheng, Yongjun He, and Jiqing Han. 2023. Graph-based spectro-temporal dependency modeling for anti-spoofing. In ICASSP.
- Chen et al. (2025) Sanyuan Chen, Chengyi Wang, Yu Wu, Ziqiang Zhang, Long Zhou, Shujie Liu, Zhuo Chen, Yanqing Liu, Huaming Wang, Jinyu Li, Lei He, Sheng Zhao, and Furu Wei. 2025. Neural Codec Language Models are Zero-Shot Text to Speech Synthesizers. IEEE Transactions on Audio, Speech and Language Processing 33 (2025), 705–718.
- Croitoru et al. (2024) Florinel-Alin Croitoru, Andrei-Iulian Hiji, Vlad Hondru, Nicolae Catalin Ristea, Paul Irofti, Marius Popescu, Cristian Rusu, Radu Tudor Ionescu, Fahad Shahbaz Khan, and Mubarak Shah. 2024. Deepfake Media Generation and Detection in the Generative AI Era: A Survey and Outlook. arXiv preprint arXiv:2411.19537 (2024).
- Frank and Schönherr (2021) J. Frank and L Schönherr. 2021. WaveFake: A Data Set to Facilitate Audio Deepfake Detection. In NeurIPS.
- Fu et al. (2021) Yihui Fu, Luyao Cheng, Shubo Lv, Yukai Jv, Yuxiang Kong, Zhuo Chen, Yanxin Hu, Lei Xie, Jian Wu, Hui Bu, et al. 2021. AISHELL-4: An Open Source Dataset for Speech Enhancement, Separation, Recognition and Speaker Diarization in Conference Scenario. In INTERSPEECH. 3665–3669.
- Gong et al. (2021) Yuan Gong, Yu-An Chung, and James Glass. 2021. AST: Audio Spectrogram Transformer. In INTERSPEECH. 571–575.
- He et al. (2016) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2016. Deep Residual Learning for Image Recognition. In CVPR. 770–778.
- Huang et al. (2022) R. Huang, M.W.Y. Lam, J. Wang, D. Su, D. Yu, Y. Ren, and Z. Zhao. 2022. FastDiff: A fast conditional diffusion model for high-quality speech synthesis. In IJCAI. 4157–4163.
- Hung et al. (2022) Yun-Ning Hung, Chih-Wei Wu, Iroro Orife, Aaron Hipple, William Wolcott, and Alexander Lerch. 2022. A Large TV Dataset for Speech and Music Activity Detection. EURASIP Journal on Audio, Speech, and Music Processing 2022, 1 (2022), 21.
- Ito and Johnson (2017) Keith Ito and Linda Johnson. 2017. The LJ Speech Dataset. https://keithito.com/LJ-Speech-Dataset/.
- Jiang et al. (2023) Ziyue Jiang, Yi Ren, Zhenhui Ye, Jinglin Liu, Chen Zhang, Qian Yang, Shengpeng Ji, Rongjie Huang, Chunfeng Wang, Xiang Yin, et al. 2023. Mega-TTS: Zero-Shot Text-to-Speech at Scale with Intrinsic Inductive Bias. arXiv preprint arXiv:2306.03509 (2023).
- Ju et al. (2024) Zeqian Ju, Yuancheng Wang, Kai Shen, Xu Tan, Detai Xin, Dongchao Yang, Eric Liu, Yichong Leng, Kaitao Song, Siliang Tang, et al. 2024. NaturalSpeech 3: Zero-Shot Speech Synthesis with Factorized Codec and Diffusion Models. In ICML.
- Jung et al. (2022) Jee-weon Jung, Hee-Soo Heo, Hemlata Tak, Hye-jin Shim, Joon Son Chung, Bong-Jin Lee, Ha-Jin Yu, and Nicholas Evans. 2022. AASIST: Audio Anti-Spoofing using Integrated Spectro-Temporal Graph Attention Networks. In ICASSP. 6367–6371.
- Kim et al. (2020) Jaehyeon Kim, Sungwon Kim, Jungil Kong, and Sungroh Yoon. 2020. Glow-tts: A generative flow for text-to-speech via monotonic alignment search. Advances in Neural Information Processing Systems 33 (2020), 8067–8077.
- Kim et al. (2021) Jaehyeon Kim, Jungil Kong, and Juhee Son. 2021. Conditional Variational Autoencoder with Adversarial Learning for End-to-End Text-to-Speech. In ICML. 5530–5540.
- Lee et al. (2023) Sang-gil Lee, Wei Ping, Boris Ginsburg, Bryan Catanzaro, and Sungroh Yoon. 2023. BigVGAN: A Universal Neural Vocoder with Large-Scale Training. In ICLR.
- Li et al. (2022) Jingyi Li, Weiping Tu, and Li Xiao. 2022. FreeVC: Towards High-Quality Text-Free One-Shot Voice Conversion. In ICASSP.
- Li et al. (2024) Yuang Li, Min Zhang, Mengxin Ren, Xiaosong Qiao, Miaomiao Ma, Daimeng Wei, and Hao Yang. 2024. Cross-Domain Audio Deepfake Detection: Dataset and Analysis. In EMNLP. 4977–4983.
- Liu et al. (2023) Xiaohui Liu, Meng Liu, Longbiao Wang, Kong Aik Lee, Hanyi Zhang, and Jianwu Dang. 2023. Leveraging positional-related local-global dependency for synthetic speech detection. In ICASSP.
- Müller et al. (2022) Nicolas M Müller, Pavel Czempin, Franziska Dieckmann, Adam Froghyar, and Konstantin Böttinger. 2022. Does audio deepfake detection generalize?. In INTERSPEECH. 2783–2787.
- Müller et al. (2024) Nicolas M Müller, Piotr Kawa, Wei Herng Choong, Edresson Casanova, Eren Gölge, Thorsten Müller, Piotr Syga, Philip Sperl, and Konstantin Böttinger. 2024. MLAAD: The Multi-Language Audio Anti-Spoofing Dataset. In IJCNN.
- Pratap et al. (2024) Vineel Pratap, Andros Tjandra, Bowen Shi, Paden Tomasello, Arun Babu, Sayani Kundu, Ali Elkahky, Zhaoheng Ni, Apoorv Vyas, Maryam Fazel-Zarandi, Alexei Baevski, Yossi Adi, Xiaohui Zhang, Wei-Ning Hsu, Alexis Conneau, and Michael Auli. 2024. Scaling Speech Technology to 1,000+ Languages. Journal of Machine Learning Research 25, 97 (2024), 1–52.
- Qin et al. (2024) Zengyi Qin, Wenliang Zhao, Xumin Yu, and Xin Sun. 2024. OpenVoice: Versatile Instant Voice Cloning. arXiv preprint arXiv:2312.01479 (2024).
- Radford et al. (2023) Alec Radford, Jong Wook Kim, Tao Xu, Greg Brockman, Christine McLeavey, and Ilya Sutskever. 2023. Robust Speech Recognition via Large-Scale Weak Supervision. In ICML. 28492–28518.
- Reimao and Tzerpos (2019) Ricardo Reimao and Vassilios Tzerpos. 2019. FoR: A dataset for synthetic speech detection. In SpeD. 1–10.
- Ristea et al. (2022) Nicolae-Cătălin Ristea, Radu Tudor Ionescu, and Fahad Shahbaz Khan. 2022. SepTr: Separable Transformer for Audio Spectrogram Processing. In INTERSPEECH. 4103–4107.
- Roselló Casado et al. (2023) Eros Roselló Casado, Alejandro Gómez Alanís, Ángel Manuel Gómez García, Antonio Miguel Peinado Herreros, et al. 2023. A conformer-based classifier for variable-length utterance processing in anti-spoofing. In INTERSPEECH. 5281–5285.
- Shen et al. (2018) Jonathan Shen, Ruoming Pang, Ron J. Weiss, Mike Schuster, Navdeep Jaitly, Zongheng Yang, Zhifeng Chen, Yu Zhang, Yuxuan Wang, Rj Skerrv-Ryan, Rif A. Saurous, Yannis Agiomvrgiannakis, and Yonghui Wu. 2018. Natural TTS Synthesis by Conditioning Wavenet on MEL Spectrogram Predictions. In ICASSP. 4779–4783.
- Shen et al. (2024) Kai Shen, Zeqian Ju, Xu Tan, Yanqing Liu, Yichong Leng, Lei He, Tao Qin, Sheng Zhao, and Jiang Bian. 2024. NaturalSpeech 2: Latent Diffusion Models are Natural and Zero-Shot Speech and Singing Synthesizers. In ICLR.
- Shi et al. (2021) Yao Shi, Hui Bu, Xin Xu, Shaoji Zhang, and Ming Li. 2021. AISHELL-3: A Multi-Speaker Mandarin TTS Corpus. In INTERSPEECH. 2756–2760.
- Solak (2019) I. Solak. 2019. The M-AILABS Speech Dataset.
- Sonobe et al. (2017) Ryosuke Sonobe, Shinnosuke Takamichi, and Hiroshi Saruwatari. 2017. JSUT corpus: free large-scale Japanese speech corpus for end-to-end speech synthesis. arXiv preprint arXiv:1711.00354 (2017).
- Suno (2023) Suno. 2023. Bark. https://github.com/suno-ai/bark
- Tak et al. (2022) Hemlata Tak, Massimiliano Todisco, Xin Wang, Jee-weon Jung, Junichi Yamagishi, and Nicholas Evans. 2022. Automatic speaker verification spoofing and deepfake detection using Wav2Vec 2.0 and data augmentation. In Odyssey. 112–119.
- Tan et al. (2024) Xu Tan, Jiawei Chen, Haohe Liu, Jian Cong, Chen Zhang, Yanqing Liu, Xi Wang, Yichong Leng, Yuanhao Yi, Lei He, Sheng Zhao, Tao Qin, Frank Soong, and Tie-Yan Liu. 2024. NaturalSpeech: End-to-End Text-to-Speech Synthesis With Human-Level Quality. IEEE Transactions on Pattern Analysis and Machine Intelligence 46, 6 (2024), 4234–4245.
- Truong et al. (2024) Duc-Tuan Truong, Ruijie Tao, Tuan Nguyen, Hieu-Thi Luong, Kong Aik Lee, and Eng Siong Chng. 2024. Temporal-Channel Modeling in Multi-head Self-Attention for Synthetic Speech Detection. In INTERSPEECH. 537–541.
- Wang et al. (2021) Changhan Wang, Morgane Riviere, Ann Lee, Anne Wu, Chaitanya Talnikar, Daniel Haziza, Mary Williamson, Juan Pino, and Emmanuel Dupoux. 2021. VoxPopuli: A Large-Scale Multilingual Speech Corpus for Representation Learning, Semi-Supervised Learning and Interpretation. In ACL. 993–1003.
- Wang et al. (2020) Xin Wang, Junichi Yamagishi, Massimiliano Todisco, Héctor Delgado, Andreas Nautsch, Nicholas Evans, Md Sahidullah, Ville Vestman, Tomi Kinnunen, Kong Aik Lee, et al. 2020. ASVspoof 2019: A large-scale public database of synthesized, converted and replayed speech. Computer Speech & Language 64 (2020), 101114.
- Yamagishi et al. (2017) Junichi Yamagishi, Christophe Veaux, and Kirsten MacDonald. 2017. CSTR VCTK Corpus: English Multi-speaker Corpus for CSTR Voice Cloning Toolkit. https://doi.org/10.7488/ds/1994 University of Edinburgh. The Centre for Speech Technology Research (CSTR).
- Yamagishi et al. (2021) Junichi Yamagishi, Xin Wang, Massimiliano Todisco, Md Sahidullah, Jose Patino, Andreas Nautsch, Xuechen Liu, Kong Aik Lee, Tomi Kinnunen, Nicholas Evans, et al. 2021. ASVspoof 2021: accelerating progress in spoofed and deepfake speech detection. In ASVspoof. 47–54.
- Yi et al. (2022) Jiangyan Yi, Ruibo Fu, Jianhua Tao, Shuai Nie, Haoxin Ma, Chenglong Wang, Tao Wang, Zhengkun Tian, Ye Bai, Cunhang Fan, et al. 2022. ADD 2022: The First Audio Deep Synthesis Detection Challenge. In ICASSP. 9216–9220.
- Yi et al. (2023) Jiangyan Yi, Jianhua Tao, Ruibo Fu, Xinrui Yan, Chenglong Wang, Tao Wang, Chu Yuan Zhang, Xiaohui Zhang, Yan Zhao, Yong Ren, et al. 2023. ADD 2023: The Second Audio Deepfake Detection Challenge. arXiv preprint arXiv:2305.13774 (2023).
- Zhang et al. (2023) Ziqiang Zhang, Long Zhou, Chengyi Wang, Sanyuan Chen, Yu Wu, Shujie Liu, Zhuo Chen, Yanqing Liu, Huaming Wang, Jinyu Li, et al. 2023. Speak foreign languages with your own voice: Cross-lingual neural codec language modeling. arXiv preprint arXiv:2303.03926 (2023).
- Zhao et al. (2023) Wenliang Zhao, Xumin Yu, and Zengyi Qin. 2023. MeloTTS: High-quality Multi-lingual Multi-accent Text-to-Speech. https://github.com/myshell-ai/MeloTTS